# MoPE: Parameter-Efficient and Scalable Multimodal Fusion via Mixture of Prompt Experts

### v2

@@ Appendix B Extended Analysis and Discussion
Ablation on vision encoder. In main experiment we use swin-base as the vision encoder. To have a fair comparision on previous prompt fusion methods (Li et al. 2023; Liang, Zhao, and Schütze 2022) that are based on ViT (Dosovitskiy et al. 2020), we provide the result of our method with ViT-base as vision encoder in Tab. 6, with results measured using 3 random seeds, As the table shows, using the ViT as vision encoder gives similar results. The result demonstrate that our superior performance is not due to more advanced vision encoder.
@@
Analysis on Adaptivity of Vanilla Prompt and MoPE
@@
Following the approach adopted in previous papers (Wang et al. 2023; Petrov, Torr, and Bibi 2023), we will focus our analysis on the case of a single prompt within a single Transformer layer. Since our method is designed to generate an effective prompt while leaving the attention calculation the same as previous prompt tuning, our MoPE will only affect the forward behavior up to the calculation of attention map. Therefore, we will concentrate on analyzing how different prompts influence the attention pattern of a pretrained Transformer layer.
@@
Theorem 1 (Limited Adaptivity of Global-shared Prompt)
@@
Furthermore, by the limited expressiveness of prompt-tuning (Petrov, Torr, and Bibi 2023), we have:
@@
Theorem 2 (Improved Adaptivity of MoPE)
@@
Thus, when the number of instances exceeds the number of experts, if the convex hull of the expert attention mappings contains an optimal attention pattern for each instance, then there exists a routing score $\mathbf{r}^{*}$ that allows MoPE to achieve an accumulated attention discrepancy equal to the sum of the optimal instance-wise attention discrepancies across those instances.
@@
Discussion: How to Choose the Main Modality?
@@
As the table shows, using BERT as the main encoder results in a decline in performance. Our postulation is that for these three tasks, the text input contains a significant amount of noisy and false positive data. For instance, in the UPMC-Food 101 dataset (Wang et al. 2015), the text data is derived using a spider, which includes many irrelevant hypertext and website titles. Therefore, treating the text as the main modality can lead to overfitting and biased predictions.
@@
Our MoPE-based method yields a better multimodal representation.
We visualize the multimodal representation generated by different methods using t-SNE. For LateConcat, this would be the concatenated feature from both modalities. For the other methods, we visualize the [CLS] token of the main modality. For ease of coloring, we only show the first 20 classes in the test set of the UPMC Food-101 (Wang et al. 2015) dataset. The results are presented in Fig. 9. As the results show, the representation generated by our method is the most separable.
@@ Appendix C Additional Visualizations
Visualization of Referring Expression Segmentation Results
In Fig. 10, Fig. 11 we provide a qualitative comparision on our method and the P-SeqFuse baseline. As illustrated in the figures, our methods could correctly understand the referring expression and localize object mask accordingly. By contrast, the compared method may fail to follow the text guidance, leading to less accurate segmentation results.
@@
Additional Examples of Expert Routing on VQAv2
Our MoPE is designed to scale up the expressiveness of vanilla prompts, and expert specialization is a critical condition to achieve superior results as well as interpretability. In the main body of the paper, we have provided visualization results of expert routing on the SNLI-VE dataset. We found the expert specialization are more observable on large dataset with a high heteriogenity. Hence, to further demonstrate expert specialization, we train our MoPE-based method on an even larger dataset, VQAv2 (Goyal et al. 2017). This dataset contains 265,016 images and paired questions, covering a wide array of visual and textual concepts. We train our model on this dataset and visualize the routing results in Figures 12, 13, 14, and 15. We observe clear expert specialization in these examples, where different experts capture different concepts.

### v3

@@ B.1 Additional Ablations
Ablation on vision encoder. In main experiment we use swin-base as the vision encoder. To have a fair comparision on previous prompt fusion methods [20, 21] that are based on ViT [6], we provide the result of our method with ViT-base as vision encoder in Tab. 7, with results measured using 3 random seeds, As the table shows, using the ViT as vision encoder gives similar results. The result demonstrate that our superior performance is not due to more advanced vision encoder.
@@
B.2 Analysis on Adaptivity of Vanilla Prompt and MoPE
@@
Following the approach adopted in previous papers [37, 28], we will focus our analysis on the case of a single prompt within a single Transformer layer. Since our method is designed to generate an effective prompt while leaving the attention calculation the same as previous prompt tuning, our MoPE will only affect the forward behavior up to the calculation of attention map. Therefore, we will concentrate on analyzing how different prompts influence the attention pattern of a pretrained Transformer layer.
@@
Theorem 1 (Limited Adaptivity of Global-shared Prompt).
@@
Furthermore, by the limited expressiveness of prompt-tuning [28], we have:
@@
Theorem 2 (Improved Adaptivity of MoPE).
@@
Thus, when the number of instances exceeds the number of experts, if the convex hull of the expert attention mappings contains an optimal attention pattern for each instance, then there exists a routing score $\mathbf{r}^{*}$ that allows MoPE to achieve an accumulated attention discrepancy equal to the sum of the optimal instance-wise attention discrepancies across those instances. ∎
@@
B.3 Discussion: How to Choose the Main Modality?
@@
As the table shows, using BERT as the main encoder results in a decline in performance. Our postulation is that for these three tasks, the text input contains a significant amount of noisy and false positive data. For instance, in the UPMC-Food 101 dataset [36], the text data is derived using a spider, which includes many irrelevant hypertext and website titles. Therefore, treating the text as the main modality can lead to overfitting and biased predictions.
@@
B.4 Our MoPE-based method yields a better multimodal representation.
We visualize the multimodal representation generated by different methods using t-SNE. For LateConcat, this would be the concatenated feature from both modalities. For the other methods, we visualize the [CLS] token of the main modality. For ease of coloring, we only show the first 20 classes in the test set of the UPMC Food-101 [36] dataset. The results are presented in Fig. 9. As the results show, the representation generated by our method is the most separable.
@@ Appendix C Additional Visualizations
C.1 Visualization of Referring Expression Segmentation Results
In Fig. 10, Fig. 11 we provide a qualitative comparison on our method and the P-SeqFuse baseline. As illustrated in the figures, our methods could correctly understand the referring expression and localize object mask accordingly. By contrast, the compared method may fail to follow the text guidance, leading to less accurate segmentation results.
@@ C.1 Visualization of Referring Expression Segmentation Results
C.2 Additional Examples of Expert Routing on VQAv2
Our MoPE is designed to scale up the expressiveness of vanilla prompts, and expert specialization is a critical condition to achieve superior results as well as interpretability. In the main body of the paper, we have provided visualization results of expert routing on the SNLI-VE dataset. We found the expert specialization are more observable on large dataset with a high heteriogenity. Hence, to further demonstrate expert specialization, we train our MoPE-based method on an even larger dataset, VQAv2 [10]. This dataset contains 265,016 images and paired questions, covering a wide array of visual and textual concepts. We train our model on this dataset and visualize the routing results in Figures 12, 13, 14, and 15. We observe clear expert specialization in these examples, where different experts capture different concepts.